let f(x)=(6/x^2 +6). find all intervals where f is concave up and all intervals where it is concave down.
1
Expert's answer
2020-08-30T17:50:25-0400

A function is concave up for the intervals where d2f(x)/dx2 >0

from f(x) =6/x2 +6, df(x)/dx = -12x-3 or df(x)/dx =-12/x3,

d2f(x)/dx2 = 36x-4 or d2f(x)/dx2 = 36/x4 +0 >0

12/x4 >0 is true.

A function f(x)=6/x2+6 is concave up for every "x \neq0."

A function is concave down for the intervals where d2f(x)/dx2 <0

from f(x) =6/x2 +6, df(x)/dx = -12x-3 or df(x)/dx =-12/x3

d2f(x)/dx2 = 36x-4 or d2f(x)/dx2 = 36/x4 +0 <0

12/x4 cannot be negative, hence the function f(x) is not concave down.